You and a friend go to a mexican restaurant. You order 2 tacos and 3 enchiladas, and your friend orders 3 tacos and 6 enchiladas. One bill is $8.00 and the other bill is $15.15. How much was each taco and each enchilada?

Answer by stanbon(75887)

You and a friend go to a mexican restaurant. You order 2 tacos and 3 enchiladas, and your friend orders 3 tacos and 6 enchiladas. One bill is $8.00 and the other bill is $15.15. How much was each taco and each enchilada?
--------
1st: 2t + 3e = 8.00
2nd: 3t + 6e = 15.15
-------------
Multiply 1st by 2 to get:
3rd: 4t + 6e = 16.00
---------
Subtract 2nd from 3rd to get:
t = 0.85 (price of a taco)
----------
Substitute that into 1st (or 2nd) to solve for e:
2*0.85 + 3e = 8.00
1.70 + 3e = 8.00
3e = 6.30
e = $2.10 (price of an enchilada)
=========
Cheers,
Stan H.